the first side of the triangle is 4 cm longer than the second side. the third side is 3 times the second side. The perimeter is 44cm. how long is each side
Found 4 solutions by Boreal, MathTherapy, ikleyn, Alan3354:
Answer by Boreal(15235)

perimeter=44=sum of sides
second is x
first is x+4
third is 3x
so x+x+4+3x=44
5x+4=44
5x=40
x=8 cm for second
12 cm for first
24 cm for third

Answer by MathTherapy(10552)

the first side of the triangle is 4 cm longer than the second side. the third side is 3 times the second side. The perimeter is 44cm. how long is each side
While solving this results in sides of 8 cm, 12 cm, and 24 cm, I must point out that such a triangle is IMPOSSIBLE to construct.
People need to do a much better job of forming these math problems. If a child who has yet to learn certain things made up such a problem,
I'd understand, but I'd also show he/she why certain NONSENSICAL problems should be avoided.
When 2 of the lines are constructed, the 3rd will either be TOO SHORT to close the triangle, or TOO LONG, thereby extending one side beyond the CLOSING point.

Answer by ikleyn(52803)

.

The tutor @Boreal correctly determined the side lengths from the condition: they are 8 cm, 12 cm and 24 cm.

But he missed to check, if such a triangle may exist.

Actually (and factually), there is NO triangle with such side lengths, since it contradicts to the triangle inequality

            a + b > c

which must be true for any combination of sides of any triangle.

In our case

            8 + 12 = 20 < 24

and the triangle inequality fails.

It is a simple rule, and I repeated it many times in my posts to this forum:

    when you solve a problem, where you need to determine the side lengths of a triangle,

    you MUST check at the end of your solution, whether the triangle inequalities are satisfied.


    Without doing it, you can make a HUGE mistake.

It is the major lesson to learn from my post.


Answer by Alan3354(69443)

this is not possible for a triangle in a plane.
I doubt it was intended to be a spherical triangle.